Liquid-nitrogen container is made of a 1 cm thick thermocoal sheet having thermal conductivity 0.025 J/m-s-°C. Liquid nitrogen at 80 k is kept in it. A total area of 0.80 m2 is in contact with the liquid nitrogen. The atmospheric temperature is 300 K. Calculate the rate of heat flow from the atmosphere to the liquid nitrogen.

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. t = 1 cm = 0.01 m, A = 0.8 m^2 θ base1 = 300, θ base 2 = 80 K = 0.025, Q/t = KA(θ base 1 – θ base 2)/l = 0.025 * 0.8 * (30030)/0.01 = 440 watt.